Consider Pacific Energy Company and U.S. Bluechips, Inc., both of which reported earnings of $957,000. Without new projects, both firms will continue to generate earnings of $957,000 in perpetuity. Assume that all earnings are paid as dividends and that both firms require a return of 14 percent. a. What is the current PE ratio for each company? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) PE ratio 7.14 times b. Pacific Energy Company has a new project that will generate additional earnings of $107,000 each year in perpetuity. Calculate the new PE ratio of the company. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) PE ratio times c. U.S. Bluechips has a new project that will increase earnings by $207,000 in perpetuity. Calculate the new PE ratio of the firm.
Answer #1
a P/E Ratio = Price / Earnings
We will use zero growth model to calculate the value of stock today.
Price = Dividend / required rate of return
Price = $957000/.14
$6835714.29
P/E ratio is calculated :
P/E = $6835714.29/$957000
PE 7.14 times
b Pacific energy company will have additional earnings of $107000 and all the earnings will
be distributed as dividend.
Total dividend paid by Pacific energy will be = $957000+ $107000
$1064000
Value of pacific energy with dividend paid $1064000
Price = $1064000/.14
$7600000
PE ratio = $7600000/$1064000
7.14 times
c U.S Bluechips will have additional earnings of $207000 and all the earnings will
be distributed as dividend.
Total dividend paid by Pacific energy will be = $957000+ $207000
$1164000
Value of pacific energy with dividend paid $1164000
Price = $1164000/.14
$8314285.71
PE ratio = $8314285.71/$1164000
7.14 times
Note: when all the earnings are distributed as dividend , The PE ratio remains
constant
Answer #2

SOLUTION :


a.


Current value of stocks of both companies 

= PV of future cash flows of $957000 in perpetuity discounted at 14%.

= 957000 / 0.14

= 6835714.29 ($)


Let number of shares be N.


So, 


Price per share = 6835714.29 / N  ($)


Earring per share = 957000 / N   ($)


PE ratio of each company.

= (6835714.29 / N) / (957000 / N) 

= 6835714.29 / 957000 

= 7.14 (ANSWER). 


b. 


With new project, earnings of Pacific energy company 

= 957000 + 107000 = 1064000 


PE ratio of this company 

= Value of company / Earnings

= (1064000 / 0.14) / 1064000 

= 7.14  (ANSWER) 


c.


With new project, earnings of US bluechip company 

= 957000 + 207000 = 1164000 


PE ratio of this company 

= Value of company / Earnings

= (1164000 / 0.14) / 1164000 

= 7.14  (ANSWER) 

answered by: Tulsiram Garg